UT 16 Sep 2021




Night Staff:

OSA: J.Williams

ISp: C. Bolyard

ISA/Observers: J.Power

Technical Wizard: J. Hill (Tucson)

Mtn Manager: P. Hartley

Sunset: 18:30

Twlight Ends: 19:48

Twilght Starts: 04:41

Sunrise: 05:59

Starting Config: MODS-Bino

Night Log:


Proposed Plan of Action:

  • MODS-Bino:
    • Spec Std at twilight
      • Dual Grating only
    • Dual BLR target - twilight until 26deg limit - end 20:40hr
    • Cataclysmic Variable - RA 20h +14 Dec - 0.5hr
    • Spectro-Std
      • All Grating Modes (if clear any seeing)
      • Skip prism modes
    • Dither test - 1.2" or better
  • If Seeing is 1.2" or better
    • LUCI-1:
      • Take LUCI-1 Rotator ZPs through Astrometric Masks to confirm no changes/issues - 0.5hr
      • New transform looks good - keep in place
    • PEPSI Training - 3 hrs
      • ready to go around 21:15 or so
      • Possible Targets:
        • RS Oph (visible until 22:10)
        • Standard stars
      • Training Availability (priority to those using in upcoming nights)
        • Setup PEPSI and Configure as if start of night
          • Andrew
          • Barry, Alex, Mark (silent mode)
  • If seeing is > 1.2"
    • PEPSI Training - 3 hrs
      • ready to go around 21:15 or so
      • Possible Targets:
        • RS Oph (visible until 22:10)
        • Standard stars
      • Training Availability (priority to those using in upcoming nights)
        • Setup PEPSI and Configure as if start of night
          • Andrew
          • Barry, Alex (silent mode)
  • After PEPSI:
    • LUCI-1 (if seeing 1.2" or better)
      • Rotator ZP Check
    • LBC:
      • If clear - standard star fields
        • Y-fan highest priority
  • Other Tasks which could be done:
    • MODS:
      • Photometric Standard Stars (if clear any seeing)
      • Pointing Model (poor weather and poor seeing)
      • NS Longslit - low priority
    • Mixed Mode Test - if time
      • LBC-Red & MODS-1
        • Test whether LBC can send a 2nd script (Dec -90) and everything works well for both sides
    • LUCI-1
      • Photometric standards
    • LBC:
      • Photometric standards (any seeing - clear)
      • Spectrophotometric standards (any seeing - clear)
  • Other Poor Weather (if pointing models are exhausted)
    • LUCI-1 local mergers
      • Service_Observing/Restart_2021/LUCI/PoorWeather/COBandhead_Mergers_V17.xml - OT file
      • finder charts in sub-directory (but they are so large you can't miss 'em)

Current Status

  • LBCs
    • LBC-Blue - available
    • LBC-Red - available
  • MODS
    • MODS-1 - issues with M1RC headers & filenumbers short 1 digit
      • AGW-5 - available
    • MODS-2 - available - glycol leak - to be inspected today
      • AGW-6 - available
  • LUCI-1 - available
    • AGW-1 available
  • LUCI-2 offline
    • AGW-2 available
  • PEPSI - available
    • AGW-7/8 available


Closed Dome Activities

  • MODS-1 Full set of Cals
    • Biases
      • 1x1
        • 8k - COMPLETED
      • 1x2
      • 2x2
    • Grating
      • Dual
        • 0.8" slit, 0.3" slit, slitless pixel flats, arcs - COMPLETED
      • Red
        • 0.8" slit, slitless pixel flats, arcs - COMPLETED
      • Blue
    • Imaging
      • Dual
      • Red
      • Blue
    • Prism
      • Dual
      • Red
      • Blue
  • MODS-2 Cals
    • Bias
      • 1x2
      • 2x2
    • Prism
      • Dual
      • Red
      • Blue
  • LUCI-1 Cals
    • Imaging N375 Flats
      • Narrow, medium-band filters
      • Broad zJHKKs and HeI completed
    • N1.8 G210 CWL=2.3 K-band 0.5" cals
    • N1.8 G200 CWL=1.93 HKspec 0.5" cals
  • ARGOS Cal-Unit test for M3 alignment and pupil wobble checkout (alternate to on-sky method)
  • LUCI-1 Only Calibrations
    • Flexure data for G150
      • /home/dthomspons/Luci/FlexureCompensation
      • README - Calib15 - 2hrs
      • Ks filter, Argon lamp, optical sieve mask
      • take dark frame
      • Manual Control MCSGUI for elevations
    • Wavelength data for G150
      • G150 Ks, Argon, take image through slit, then put in grating, line should show up where the slit is located (within a few pixels of the image of the slit)
      • Check a few lines (2 or 3)

Night Overview

LUCI, MODS, and LBC's started up

25 LBC biases taken, LUCI field stop verified

MODS Skyflats

1:50 Sending test to blank 17+66 for mods skyflats for dual imaging with g & r

2:05 We are calling it with skyflats, past threshold for off axis guiding and want to move to first field for pointing and collimation check. MODS1 blue flat times were not great (I struggled to update that one on time) others were ok). Obtained a g&3r set.
Slewing to BD33

2:13 Pointing has been checking and sending acquisition script

MODS1 Offset Command:
offsetxy -0.926 12.400 rel

MODS2 Offset Command:
offsetxy 3.578 9.396 rel

SpecPhot

2:19 Starting BD33 science

mods1b.20210916.0022-24
mods1r.20210916.0021-23
mods2b.20210916.0021-23
mods2r.20210916.0021-23
Seeing 0.78 on Dimm, 0.8-0.9" on the guider

Dual BLR

2:35 Going to Dual BLR

Computed Slit Alignment Offset:
dX = -0.858 arcsec
dY = 12.656 arcsec

MODS1 Offset Command:
offsetxy -0.858 12.656 rel

MODS2 Offset Command:
offsetxy 3.471 10.127 rel

2:46 Science!

We can see the LED we saw yesterday. Chris is heading up to take a look. Elevation is 34.  


LED.png:  

LED.png

Stop and readout for Dual BLR target. Had to have josh stop the drives because mods readout lag was just a little too long. Dang.

j2055

3:47 Slewing to j2055. 0.6" on the guider, clear. Updating script to use the 0.8" slit.

4:07 Starting science! 0.65-0.75" on the guider. Clear conditions.

mods1b.20210916.0027-37
mods1r.20210916.0034-44
mods2b.20210916.0026-36
mods2r.20210916.0034-44

Dither Test

4:55 slewing for MODS dither test on M52. Running M52_pointing_abs_newGS. Once again GCS looks a bit cut off on the top on the left and right. IT looked well centered on our previous target. Having Josh restart GCS before we proceed with the dither test to see if this is recsolved.

5:00 Resending preset and starting. Initial offsets to align stars were -5.8 on MODS1 and -6.8 on MODS2. The Guider WFS shift persisted after the restart of GCS. Seeing is 0.8" on teh guiders and 1" on the DIMM. We seem well collimated,
WFS.png:
WFS.png

Spectrophotometric Standard in Dual and Red Grating

5:54 Homing guide probe before slewing to see if that helps with WFS lenslet vignette. Sending preset to Feige 110

MODS1 Offset Command:
offsetxy -0.853 11.652 rel

MODS2 Offset Command:
offsetxy 3.613 8.671 rel, additional -0.25" needed in X

Seeing 0.57" on the DIMM 0.65" on the guider.

6:06 Starting science for dual grating.

mods1b.20210916.0068-70
mods1r.20210916.0077-79
mods2b.20210916.0057-59
mods2r.20210916.0078-80

6:17 Starting Red Grating Science, seeing 0.59" on the dimm and 0.7" on the guider.

mods1r.20210916.0080
mods2r.20210916.0081

LUCI Rot ZP verification

6:29 Slewing up to reconfigure to LUCI1

6:47 Sending preset for RotZP check on stnP2020 mask. Seeing is 0.6" on the guider and WFS,

Removed star 36. It's bright and likely had proper motions that needed to be taken into account from last years generation.
X-Shift -0.33176
Y-Shift -0.45282
PA 0.07701

Not quite... Stars in the lower right quadrant appear shifted down and to the right. Is it the solution or is it the field. Going to do another field to verify.

7:01 Sewing to stoneA for a second verification. No stars removed from Stone A. Seeing 0.76" on WFS. 0.83" on DIMM
X-Shift -0.91660
Y-Shift -0.53300
PA 0.06858

Astrometric alignment:
luci_stnA2020.png

Confirmation image: all of the stars are sfited to the right: 0.45". Offsetting and taking an additional confirmation image. Does not look like a rotation. Looks perfect!!

Looks like the Stone P 2020 mask may contain a number of bright stars and may be plagues by proper motions. Should be removed from astrometric rotation

7:19 Reconfiguring to PEPSI

7:47 After some fun trying to find trying to find where the new pointing model was located and updating it, updated PEPSI pointing model on /lbt/data/config/tcs/pcs SX_PEPSIv1 and DX_PEPSIv1. Apparently the two files cannot be binocularized. Need to be edited to remove

7:52 Restarting PCS agian with fixed mocularized pointing files. Sending acquire preset. first preset was in teh FoV so thats an improvement. We will leave this pointing

8:00 Starting PEPSI training.

Sorry for the lack of details, but training went well. Pointing was bang on. We did a series of pointings around the sky and wrapped up

9:50 Andrew has turned on the TMS laser. We are wrapping up with a flux standard with PEPSI. Seeing 1" on the DIMM, 0.75" on teh guider and photmetric.

16/09/2021 08:01:28.7 00:02:59.301 BD+28 4211 3: 4800 - 5441 300 194 0.74 pepsib.20210916.000

16/09/2021 08:01:28.7 00:02:59.299 BD+28 4211 5: 6278 - 7419 300 175 0.74 pepsir.20210916.000

16/09/2021 08:29:04.7 00:01:00.000 beta Cep 3: 4800 - 5441 300 2659 0.94 pepsib.20210916.001

16/09/2021 08:29:04.7 00:01:00.000 beta Cep 5: 6278 - 7419 300 2520 0.94 pepsir.20210916.001

16/09/2021 08:37:27.2 00:04:00.000 HIP 5034 5: 6278 - 7419 300 1730 0.76 pepsir.20210916.002

16/09/2021 08:37:27.2 00:04:00.000 HIP 5034 3: 4800 - 5441 300 1436 0.76 pepsib.20210916.002

16/09/2021 08:47:11.8 00:05:00.000 HIP 9780 3: 4800 - 5441 300 857 1.73 pepsib.20210916.003

16/09/2021 08:47:11.8 00:05:00.000 HIP 9780 5: 6278 - 7419 300 1730 1.73 pepsir.20210916.003

16/09/2021 09:01:00.9 00:01:00.000 beta Cep 5: 6278 - 7419 300 3873 1.09 pepsir.20210916.004

16/09/2021 09:01:00.9 00:01:00.000 beta Cep 3: 4800 - 5441 300 4141 1.09 pepsib.20210916.004

16/09/2021 09:11:26.6 00:05:00.000 VX Cas 5: 6278 - 7419 300 162 0.59 pepsir.20210916.005

16/09/2021 09:11:26.6 00:05:00.000 VX Cas 3: 4800 - 5441 300 131 0.59 pepsib.20210916.005

16/09/2021 09:20:34.8 00:05:00.000 HD 17156 5: 6278 - 7419 300 0.82 pepsir.20210916.006

16/09/2021 09:20:34.8 00:05:00.000 HD 17156 3: 4800 - 5441 300 0.82 pepsib.20210916.006

16/09/2021 09:29:28.3 00:01:00.000 phi Per 3: 4800 - 5441 300 0.70 pepsib.20210916.007

16/09/2021 09:29:28.2 00:01:00.000 phi Per 5: 6278 - 7419 300 0.70 pepsir.20210916.007

16/09/2021 09:37:02.7 00:05:00.000 HIP 454 5: 6278 - 7419 300 0.80 pepsir.20210916.008

16/09/2021 09:37:02.7 00:05:00.000 HIP 454 3: 4800 - 5441 300 0.80 pepsib.20210916.008

16/09/2021 09:45:11.2 00:01:00.000 HR 9087 3: 4800 - 5441 100 239 fluxstd 0.82 pepsib.20210916.009

16/09/2021 09:45:11.2 00:01:00.000 HR 9087 5: 6278 - 7419 100 167 fluxstd 0.82 pepsir.20210916.009

16/09/2021 09:51:09.4 00:01:00.000 HR 9087 2: 4265 - 4800 100 124 fluxstd 0.74 pepsib.20210916.010

16/09/2021 09:51:09.4 00:01:00.000 HR 9087 4: 5441 - 6278 100 204 fluxstd 0.74 pepsir.20210916.010

16/09/2021 09:53:38.4 00:01:00.000 HR 9087 6: 7419 - 9067 100 172 fluxstd 0.73 pepsir.20210916.011

16/09/2021 09:53:38.4 00:01:00.000 HR 9087 1: 3837 - 4265 100 121 fluxstd 0.73 pepsib.20210916.011

16/09/2021 09:57:24.4 00:01:00.000 HR 9087 5: 6278 - 7419 200 578 fluxstd 0.74 pepsir.20210916.012

16/09/2021 09:57:24.4 00:01:00.000 HR 9087 3: 4800 - 5441 200 606 fluxstd 0.74 pepsib.20210916.012

16/09/2021 09:59:11.8 00:01:00.000 HR 9087 2: 4265 - 4800 200 522 fluxstd 0.72 pepsib.20210916.013

16/09/2021 09:59:11.8 00:01:00.000 HR 9087 4: 5441 - 6278 200 545 fluxstd 0.72 pepsir.20210916.013

16/09/2021 10:01:01.6 00:01:00.000 HR 9087 6: 7419 - 9067 200 488 fluxstd 0.74 pepsir.20210916.014

16/09/2021 10:01:01.6 00:01:00.000 HR 9087 1: 3837 - 4265 200 277 fluxstd 0.74 pepsib.20210916.014

16/09/2021 10:04:22.6 00:01:00.000 HR 9087 3: 4800 - 5441 300 1088 fluxstd 0.79 pepsib.20210916.015

16/09/2021 10:04:22.6 00:01:00.000 HR 9087 5: 6278 - 7419 300 1144 fluxstd 0.79 pepsir.20210916.015

16/09/2021 10:05:56.5 00:01:00.000 HR 9087 4: 5441 - 6278 300 1130 fluxstd 0.81 pepsir.20210916.016

16/09/2021 10:05:56.5 00:01:00.000 HR 9087 2: 4265 - 4800 300 957 fluxstd 0.81 pepsib.20210916.016

16/09/2021 10:07:42.7 00:01:00.000 HR 9087 6: 7419 - 9067 300 1023 fluxstd 0.80 pepsir.20210916.017

16/09/2021 10:07:42.7 00:01:00.000 HR 9087 1: 3837 - 4265 300 650 fluxstd 0.80 pepsib.20210916.017

10:09 Reconfiguring for binocular LBC's

10:33 Sending preset to First LBC phot field Y5. TMS channel 10 & 21 are dropped. It looks like some light on the chamber floor
  • Primary Cam light.png:
    Primary Cam light.png

10:40 LBCRed first dohybrid.png. Something is funky with the LBCR images when running dohybrid. Stray light? Josh is on the hunt.
LBCRed first dohybrid.png

10:43 Josh is awesome and found the source. Lights left on on 4. All good now
Stray light fixed.png:
Stray light fixed.png

8:48 sending copointing. Copointing failed with persistent iraf error. But we look well copointed. So we will skip this for now.

8:57 Copointing skipped. TMS setpoint set and we are starting science on sdssY5_r0330d04

image 110422 looks like we had a small image jump

111005 looks really out of focus, there was a large shift in TMS reported by John.

Looks like we may be having the filter problem with TMS

11:29 Sending to sdssY7. Restarted iraf so will try the copointing again at this lower elevation. Will not collimate, trusting active TSM to handle the fun stuff. John fixed a line in the code to help with the timeing for filters and TMS

11:34 Copointing completed. Starting science. There is a huge correction in huge group rotation and the images do not look good. So lesson of the day is that a new reference needs to be taken after copointing. Does TMS undo copointing? So what does this mean for yesterdays pointing model. We did take long 20sec images after several pointing model copointed images yesterday. We should check to see if TMS uncopointed those. Rerunning dofpia and copointing. Resetting setpoint after copoint

11:48 sending sdssY7

115925 too few downloaded pixels

The last images are definitely impacted by sky brightness.

TMS jump on 121417, likely because of filter offset.

Channels 21 and 10 are both giving no signal at elevation 48.

12:24 TMS lasers off

LBC Skyflats

12:22 Skyflat V & r at PA0 & PA180

12:37 SkyFlats B &z at PA180 , too bright for PA0

Calibrations

LBC 25 biases

MODS: 1x2 bias, 1x2 dual grating arcs,

Pausing cals for Chris to do fills and mask exchange.

1x2 0.8" dual grating longslit flats, 1x2 dual grating pixel flats

red grating arcs, red grating slitless, red grating 5" long slit

dual grating arcs, dual grating flats 5"

-- BarryRothberg - 15 Sep 2021

*
I Attachment Action Size Date Who Comment
LBCRed first dohybrid.pngpng LBCRed first dohybrid.png manage 1 MB 16 Sep 2021 - 10:40 JenniferPower  
LED.pngpng LED.png manage 1 MB 16 Sep 2021 - 02:55 JenniferPower  
Primary Cam light.pngpng Primary Cam light.png manage 2 MB 16 Sep 2021 - 10:38 JenniferPower  
Stray light fixed.pngpng Stray light fixed.png manage 1 MB 16 Sep 2021 - 10:47 JenniferPower  
WFS.pngpng WFS.png manage 639 K 16 Sep 2021 - 05:17 JenniferPower  
luci_stnA2020.pngpng luci_stnA2020.png manage 104 K 16 Sep 2021 - 12:55 JenniferPower Astrometric alignment
Topic revision: r7 - 16 Sep 2021, JenniferPower
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback